Suppose the demand curve for a product is given by Q = 300 − 2P + 4I, where I is average income measured in thousands of dollars. The supply curve is Q = 3P − 50.

If I = 25, find the market-clearing price and quantity for the product.

If I = 50, find the market-clearing price and quantity for the product.

Draw a graph to illustrate your answers.
1
Expert's answer
2020-12-10T14:29:33-0500

If I = 25

"Q = 300 - 2P + 4 \times 25"

Demand curve

Q = 400 - 2P

Supply curve = Q = 3P - 50

Frr market price and quantity, demand = supply.

400 - 2P = 3P - 50

5P = 450

P = 90

400 - 180 = 220

Therefore, market price is 90 and quantity is 220.

If I = 50

demand curve "= 300 - 2P + 4 \times 50 = 500 - 2P"

supply = 3P - 50

500 - 2P = 3P -50

5P = 550

P = 110

Q = 500 - 220 = 280

The market price is 110 and quantity is 280.